A cord passing over a frictionless, massless pulley has a 4 kg. object tied to one end and 12 kg object tied to the other. compute the acceleration and the tension in the cord.


1
Expert's answer
2021-05-11T09:07:07-0400

(a) Let's apply the Newton's Second Law of Motion:


"T-m_1g=m_1a,""T-m_2g=-m_2a."

Eliminating "T", we get:


"(m_2-m_1)g=(m_1+m_2)a,""a=\dfrac{(m_2-m_1)g}{m_1+m_2},""a=\dfrac{(12\ kg-4\ kg)\cdot9.8\ \dfrac{m}{s^2}}{4\ kg+12\ kg}=4.9 \ \dfrac{m}{s^2}."

(b) We can find the tension in the cord by substituting "a" into the first equation:


"T=m_1(a+g)=4\ kg\cdot(4.9 \ \dfrac{m}{s^2}+9.8 \ \dfrac{m}{s^2})=58.8\ N."